OpenStudy (haleyelizabeth2017):

Find a quadratic equation with roots -1+4i and -1-4i @TheSmartOne

TheSmartOne (thesmartone):

Well If the solution is 1 and 2 Then it is (x-1)(x-2)=0 So for this one can you guess from here?

OpenStudy (haleyelizabeth2017):

Isn't there only supposed to be like (x-_)(x-_) because if I solve it like that, then there is gonna be (x-_+_)(x-_-_).......so I am completely lost on this :(

TheSmartOne (thesmartone):

Yes.

TheSmartOne (thesmartone):

So \[(x+(1-4i))(x+(1+4i))\]

OpenStudy (haleyelizabeth2017):

oh....okay.....

OpenStudy (haleyelizabeth2017):

so do I just do something like FOIL to find the equation?

TheSmartOne (thesmartone):

Yes...

TheSmartOne (thesmartone):

And remember\[i^2=\sqrt{-1}\]

OpenStudy (campbell_st):

you could use \[x = -1 \pm \sqrt{-16}\] and work backwards

OpenStudy (haleyelizabeth2017):

\(i^2\) is equal to -1 not \[\sqrt{-1}\] I thought......:(

TheSmartOne (thesmartone):

oops yes.

OpenStudy (campbell_st):

then \[(x + 1) = \pm \sqrt{-16}\] square both sides

OpenStudy (haleyelizabeth2017):

Okay.....sorry @campbell_st! I think I'm gonna stick with TheSmartOne's method....thank you though

OpenStudy (haleyelizabeth2017):

so anyhoo.....how do I do it when (1-4i) is in parenthesis? I am so sorry but I am not getting math very much tonight....:(

TheSmartOne (thesmartone):

\[(x+1-4i)(x+1+4i)=0\]

OpenStudy (campbell_st):

ok... if you square both sides you get \[(x + 1)^2 = -16\] all you need to know is that \[\pm 4i = \pm \sqrt{-16}\] good luck

OpenStudy (haleyelizabeth2017):

so is it just like FOIL but with 3 different ones instead of two?

OpenStudy (campbell_st):

yes... and your diffuclty will be the expansion.

TheSmartOne (thesmartone):

We would get \[x^2+x+4ix+x+1+4i-4ix-4i-4i^2=0\]

OpenStudy (haleyelizabeth2017):

okay thats what I got on paper.....then simplify to \(x^2+2x+1-36i\) because the +4ix and -4ix cancel out......

OpenStudy (haleyelizabeth2017):

whoops! I got that wrong......

TheSmartOne (thesmartone):

Oh oops.

OpenStudy (haleyelizabeth2017):

so \(x^2+2x+1-20i-16i^2\)

TheSmartOne (thesmartone):

\(x^2+x+4ix+x+1+4i-4ix-4i-16i^2=x^2+2x+1+8i-16i^2\)

OpenStudy (haleyelizabeth2017):

so that would actually be \(x^2+2x+17-20i\) because -16*-1=+16 and that can combine (like terms) with +1

OpenStudy (haleyelizabeth2017):

I am unsure what -20i would turn into so we could simplify it even further....:(

TheSmartOne (thesmartone):

How did you get -20i ??

OpenStudy (haleyelizabeth2017):

never mind......my mistake ha ha yeah......:(

OpenStudy (haleyelizabeth2017):

sorry

TheSmartOne (thesmartone):

my mistake too -4i+4i means no i

TheSmartOne (thesmartone):

Let me go back and show you.. I made a few mistakes lol

OpenStudy (haleyelizabeth2017):

so it is just \(x^2+2x+17\) :D

TheSmartOne (thesmartone):

\[(x+1-4i)(x+1+4i)=0\] \[x^2+1x+4ix+1x+1+4i-4ix-4i-16i^2=0\] \[x^2+4ix-4ix+1x+1x+4i-4i-16i^2=0\] \[x^2+2x-16i^2=0\]

TheSmartOne (thesmartone):

\(\ i^2=-1\) So \(\ x^2+2x-16i^2=x^2+2x-16(-1)\) \(\ x^2+2x+16\)

OpenStudy (haleyelizabeth2017):

+16? what happened to the +1?

TheSmartOne (thesmartone):

Oops all these errors..

TheSmartOne (thesmartone):

Lets backtrack.

OpenStudy (haleyelizabeth2017):

haha I got it though.....you just forgot to add the 1 so it should be \(x^2+2x+17\) :D

TheSmartOne (thesmartone):

\(\ x^2+1x+4ix+1x+1+4i-4ix-4i-16i^2=0\) \(\ x^2+4ix-4ix+1x+1x+4i-4i-16i^2+1=0\) \(\ x^2+2x-16i^2+1=0\) \(\ x^2+2x-16i^2+1=x^2+2x-16(-1)+1\) \(\ x^2+2x+17\)

TheSmartOne (thesmartone):

Finally

OpenStudy (haleyelizabeth2017):

Thank you sooooo sooooo much for all the help :D you have NO idea how much I appreciate your help :) thanks again

TheSmartOne (thesmartone):

No problem... Sorry for all the mistakes lol

OpenStudy (haleyelizabeth2017):

hey, it's okay! We got the answer in the end still! :D

TheSmartOne (thesmartone):

:)